An iron railway line is 30m long at a temperature of 27oC. What would be the length of the 

railway line when the temperature increases to 57oC? The coefficient of linear expansion of 

iron is 1.2 x 10-5 oC-1



1
Expert's answer
2021-10-02T15:44:23-0400

Solution.

"L=30m;"

"t_1=27^oC;"

"t_2=57^oC;"

"\alpha=1.2\sdot10^{-5}C^{-1};"

"L_0=\dfrac{L}{1+\alpha \Delta T};"

"L_0=\dfrac{30}{1+1.2\sdot10^{-5}\sdot7}=29.99m;"

"L1=L_0(1+\alpha \Delta T);"

"L_1=29.99(1+1.2\sdot10^{-5}\sdot37)=30.003m;"

Answer: "L_1=30.003m."